THE MAGELLANIC CLOUDS NEWSLETTER

An electronic publication dedicated to the Magellanic Clouds, and astrophysical phenomena therein

No. 88 — 1 August 2007

http://www.astro.keele.ac.uk/MCnews

Editors: Jacco van Loon and Snežana Stanimirović

Editorial

Dear Colleagues,

It is our pleasure to present the 88^{th} issue of the Magellanic Clouds Newsletter. With 19 abstracts of refereed journal papers and two postgraduate positions opening to work in the UK and Italy specifically on the Magellanic Clouds, an ever-growing number of registrants to the newsletter and the Sydney workshop on the Magellanic Clouds just behind us, we can look forward to a bright future of Magellanic Clouds research. Topics that are especially well represented in the current issue of the newsletter are star clusters, Cepheid variables, X-ray binaries and the structure of the Magellanic System.

The 2007 Magellanic Clouds observing season is starting in earnest now, for the groundbased optical and infrared astronomers amongst us. We are curious to know how early or late in the season people have still be able to take (publishable) data of the Magellanic Clouds. Do let us know, if you observed the SMC in March or April, the LMC in June, or if you managed to obtain useful low-resolution optical spectra by imaging close to the horizon.

The next issue will be distributed on the 2^{nd} of October; the deadline for contributions is the 1^{st} of October.

Editorially Yours,

Jacco van Loon and Snežana Stanimirović

Refereed Journal Papers

Surface Brightness Profiles for a sample of LMC, SMC and Fornax galaxy Globular Clusters

Eva Noyola^{1,2} and Karl Gebhardt^{1}

1University of Texas at Austin, USA 2Max-Planck-Institut für extraterrestrische Physik, Germany

We use Hubble Space Telescope archival images to measure central surface brightness profiles of globular clusters around satellite galaxies of the Milky Way. We report results for 21 clusters around the LMC, 5 around the SMC, and 4 around the Fornax dwarf galaxy. The profiles are obtained using a recently developed technique based on measuring integrated light, which is tested on an extensive simulated dataset. Our results show that for 70% of the sample, the central photometric points of our profiles are brighter than previous measurements using star counts with deviations as large as 2 mag acrsec^{−2}. About 40% of the objects have central profiles deviating from a flat central core, with central logarithmic slopes continuously distributed between −0.2 and −1.2. These results are compared with those found for a sample of Galactic clusters using the same method. We confirm the known correlation in which younger clusters tend to have smaller core radii, and we find that they also have brighter central surface brightness values. This seems to indicate that globular clusters might be born relatively concentrated, and that a profile with extended flat cores might not be the ideal choice for initial profiles in theoretical models.

Accepted for publication in Astronomical Journal Available from arXiv:0705.3464

Clustered Star Formation in the Small Magellanic Cloud. A Spitzer/IRAC View of the Star-Forming Region NGC 602/N 90

Dimitrios A. Gouliermis^{1}, Sascha P. Quanz^{1} and Thomas Henning^{1}

1Max-Planck-Institut für Astronomie, Königstuhl 17, 69117 Heidelberg, Germany

We present Spitzer/IRAC photometry on the star-forming H ii region N 90, related to the young stellar association NGC 602 in the Small Magellanic Cloud. Our photometry revealed bright mid-infrared sources, which we classify with the use of a scheme based on templates and models of red sources in the Milky Way, and criteria recently developed from the Spitzer Survey of the SMC (Bolatto et al. 2007) for the selection of candidate Young Stellar Objects (YSOs). We detected 57 sources in all four IRAC channels in a 6.2′ × 4.8′ field-of-view centered on N 90; 22 of these sources are classified as candidate YSOs. We compare the locations of these objects with the position of optical sources recently found in the same region with high-resolution HST/ACS imaging of NGC 602 by Schmalzl et al. (2007), and we find that 17 candidate YSOs have one or more optical counterparts. All of these optical sources are identified as pre-main sequence stars, indicating, thus, ongoing clustered star formation events in the region. The positions of the detected YSOs and their related PMS clusters give a clear picture of the current star formation in N 90, according to which the young stellar association photo-ionizes the surrounding interstellar medium, revealing the H ii nebula, and triggering sequential star formation events mainly along the eastern and southern rims of the formed cavity of the parental molecular cloud.

Accepted for publication in Astrophysical Journal Available from arXiv:0705.3359

Discovery of a nearby twin of SN 1987A’s nebula around the luminous blue variable HD 168625: Was Sk–69 202 an LBV?

Nathan Smith^{1}

1Astronomy Department, University of California, 601 Campbell Hall, Berkeley, CA 94720, USA

Spitzer images of the luminous blue variable (LBV) candidate HD 168625 reveal the existence of a bipolar nebula several times larger than its previously-known equatorial dust torus. The outer nebula of HD 168625 has a full extent of 80′′ or 0.85 pc, and one of the lobes has a well-defined polar ring. The nebula is a near twin of the triple-ring system around SN 1987A. Because of these polar rings, and accounting for stellar/progenitor luminosity, HD 168625 is an even closer twin of SN 1987A than the B supergiant Sher 25 in NGC 3603. HD 168625’s nebula was probably ejected during a giant LBV eruption and not during a red supergiant phase, so its similarity to the nebula around SN 1987A may open new possibilities for the creation of SN1987A’s rings. Namely, the hypothesis that Sk–69 202 suffered an LBV-like eruption would avert the complete surrender of single star models for its bipolar nebula by offering an alternative to an unlikely binary merger scenario. It also hints that LBVs are the likely progenitors of some type II supernovae, and that HD 168625’s nebula is a good example of a pre-explosion environment.

Published in AJ, 133, 1034 (2007) Available from astro-ph/0611544

Semi-Empirical Cepheid Period-Luminosity Relations in Sloan Magnitudes

C. Ngeow^{1} and S. Kanbur^{2}

1UIUC, USA 2SUNY-Oswego, USA

In this paper we derive semi-empirical Cepheid period-luminosity (P-L) relations in the Sloan ugriz magnitudes by combining the observed BVI mean magnitudes from the Large Magellanic Cloud Cepheids (LMC) and theoretical bolometric corrections. We also constructed empirical gr band P-L relations, using the publicly available Johnson-Sloan photometric transformations, to be compared with our semi-empirical P-L relations. These two sets of P-L relations are consistent with each other.

Accepted for publication in ApJ Available from arXiv:0706.1537

The Detailed Forms of the LMC Cepheid PL and PLC Relations

C. Koen^{1}, S. Kanbur^{2} and C. Ngeow^{3}

1University of the Western Cape, South Africa 2SUNY-Oswego, USA 3UIUC, USA

Possible deviations from linearity of the LMC Cepheid PL and PLC relations are investigated. Two datasets are studied, respectively from the OGLE and MACHO projects. A nonparametric test, based on linear regression residuals, suggests that neither PL relation is linear. If colour dependence is allowed for then the MACHO PL relation is found to deviate more significantly from the linear, while the OGLE PL relation is consistent with linearity. These finding are confirmed by fitting “Generalised Additive Models” (nonparametric regression functions) to the two datasets. Colour dependence is shown to be nonlinear in both datasets, distinctly so in the case of the MACHO Cepheids. It is also shown that there is interaction between the period and colour functions in the MACHO data.

Accepted for publication in MNRAS Available from arXiv:0706.1762

Period-colour and amplitude-colour relations in classical Cepheid variables V: The Small Magellanic Cloud Cepheid models

S. Kanbur^{1}, C. Ngeow^{2} and G. Feiden^{1}

1SUNY-Oswego, USA 2UIUC, USA

Period-colour (PC) and amplitude-colour (AC) relations at maximum, mean and minimum light are constructed from a large grid of full amplitude hydrodynamic models of Cepheids with a composition appropriate for the SMC (Small Magellanic Cloud). We compare these theoretical relations with those from observations. The theoretical relations are in general good agreement with their observational counterparts though there exist some discrepancy for short period (log[P] < 1) Cepheids. We outline a physical mechanism which can, in principle, be one factor to explain the observed PC/AC relations for the long and short period Cepheids in the Galaxy, LMC and SMC. Our explanation relies on the hydrogen ionization front-photosphere interaction and the way this interaction changes with pulsation period, pulsation phase and metallicity. Since the PC relation is connected with the period-luminosity (PL) relation, it is postulated that such a mechanism can also explain the observed properties of the PL relation in these three galaxies.

Accepted for publication in MNRAS Available from arXiv:0706.2629

Chemical evolution of the Small Magellanic Cloud based on planetary nebulae

T. P. Idiart^{1}, W. J. Maciel^{1} and R. D. D. Costa^{1}

1IAG/USP, Rua do Matão 1226, CEP 05508-900, São Paulo SP, Brazil

We investigate the chemical evolution of the Small Magellanic Cloud (SMC) based on abundance data of planetary nebulae (PNe). The main goal is to investigate the time evolution of the oxygen abundance in this galaxy by deriving an age-metallicity relation. Such a relation is of fundamental importance as an observational constraint of chemical evolution models of the SMC. We have used high quality PNe data in order to derive the properties of the progenitor stars, so that the stellar ages could be estimated. We collected a large number of measured spectral fluxes for each nebula, and derived accurate physical parameters and nebular abundances. New spectral data for a sample of SMC PNe obtained between 1999 and 2002 are also presented. These data are used together with data available in the literature to improve the accuracy of the fluxes for each spectral line. We obtained accurate chemical abundances for PNe in the Small Magellanic Cloud, which can be useful as tools in the study of the chemical evolution of this galaxy and of Local Group galaxies. We present the resulting oxygen versus age diagram and a similar relation involving the [Fe/H] metallicity based on a correlation with stellar data. We discuss the implications of the derived age-metallicity relation for the SMC formation, in particular by suggesting a star formation burst in the last 2-3 Gyr.

Accepted for publication in Astronomy and Astrophysics Available from arXiv:0706.1967 and from http://www.astro.iag.usp.br/∼maciel

On the Interpretation of the Age Distribution of Star Clusters in the Small Magellanic Cloud

M. Gieles^{1,2}, H.J.G.L.M. Lamers^{1} and S.F. Portegies Zwart^{3}

1Utrecht University, The Netherlands 2ESO/Santiago, Chile 3University of Amsterdam, The Netherlands

We re-analyze the age distribution (dN/dt) of star clusters in the Small Magellanic Cloud (SMC) using age determinations based on the Magellanic Cloud Photometric Survey. For ages younger than 3 × 10^{9} yr the dN/dt distribution can be approximated by a power-law distribution, dN/dt ∝ t^{−β}, with −β = −0.70 ± 0.05 or −β = −0.84 ± 0.04, depending on the model used to derive the ages. Predictions for a cluster population without dissolution limited by a V-band detection result in a power-law dN/dt distribution with an index of ∼ −0.7. This is because the limiting cluster mass increases with age, due to evolutionary fading of clusters, reducing the number of observed clusters at old ages. When a mass cut well above the limiting cluster mass is applied, the dN/dt distribution is flat up to 1 Gyr. We conclude that cluster dissolution is of small importance in shaping the dN/dt distribution and incompleteness causes dN/dt to decline. The reason that no (mass independent) infant mortality of star clusters in the first ∼ 10-20 Myr is found is explained by a detection bias towards clusters without nebular emission, i.e. cluster that have survived the infant mortality phase. The reason we find no evidence for tidal (mass dependent) cluster dissolution in the first Gyr is explained by the weak tidal field of the SMC. Our results are in sharp contrast to the interpretation of Chandar et al. (2006), who interpret the declining dN/dt distribution as rapid cluster dissolution. This is due to their erroneous assumption that the sample is limited by cluster mass, rather than luminosity.

Accepted for publication in Astrophysical Journal Available from arXiv:0706.1202

Be stars and binaries in the field of the SMC open cluster NGC 330 with VLT-FLAMES

C. Martayan ^{1}, M. Floquet ^{1}, A.M. Hubert ^{1}, J. Gutiérrez-Soto ^{1,2}, J. Fabregat ^{2}, C. Neiner ^{1} and M. Mekkas ^{1}

1GEPI, Observatoire de Paris, CNRS, Université Paris Diderot; place Jules Janssen 92195 Meudon Cedex, France 2Observatorio Astronómico de Valencia, edifici Instituts d’investigació, Poligon la Coma, 46980 Paterna Valencia, Spain

Observations of hot stars belonging to the young cluster SMC-NGC 330 and its surrounding region were obtained with the VLT-GIRAFFE facilities in MEDUSA mode. We investigated the B and Be star properties and proportions in this environment of low metallicity. We also searched for rapid variability in Be stars using photometric databases. With spectroscopic measurements we characterized the emission and properties of Be stars. By cross-correlation with photometric databases such as MACHO and OGLE, we searched for binaries in our sample of hot stars, as well as for short-term variability in Be stars. We report on the global characteristics of the Be star sample (131 objects). We find that the proportion of early Be stars with a large equivalent width of the Hα emission line is higher in the SMC than in the LMC and MW. We find a slight increase in the proportion of Be stars compared to B-type stars with decreasing metallicity. We also discovered spectroscopic and photometric binaries, and for the latter we give their orbital period. We identify 13 Be stars with short-term variability. We determine their period(s) and find that 9 Be stars are multiperiodic.

Accepted for publication in A&A Available from arXiv:0706.4426

Stellar Populations in the Outskirts of the Small Magellanic Cloud: No Outer Edge Yet

Noelia E. D. Noël^{1} and Carme Gallart^{1}

1Instituto de Astrofísica de Canarias, Spain

We report the detection of intermediate-age and old stars belonging to the SMC at 6.5 kpc from the SMC center in the southern direction. We show, from the analysis of three high quality 34′ × 33′ CMDs, that the age composition of the stellar population is similar at galactocentric distances of ∼ 4.7 kpc, ∼ 5.6 kpc, and ∼ 6.5 kpc. The surface brightness profile of the SMC follows an exponential law, with no evidence of truncation, all the way out to 6.5 kpc. These results, taken together, suggest that the SMC ‘disk’ population is dominating over a possible old Milky Way-like stellar halo, and that the SMC may be significantly larger than previously thought.

Accepted for publication in ApJ Letters Available from arXiv:0707.0312

Discovery of the Pre-Main Sequence Population of the Stellar Association LH 95 in the Large Magellanic Cloud with Hubble Space

Telescope ACS Observations

Dimitrios A. Gouliermis^{1}, Thomas Henning^{1}, Wolfgang Brandner^{1,2}, Andrew E. Dolphin^{3}, Michael Rosa^{4,5} and Bernhard Brandl^{6}

1Max Planck Institute for Astronomy, Heidelberg, Germany 2UCLA, Div. of Astronomy, Los Angeles, USA 3Raytheon Corporation, USA 4Space Telescope European Coordinating Facility, ESO, Germany 5Space Telescope Operations Division, RSSD, ESA 6Leiden University, Leiden Observatory, The Netherlands

We report the discovery of an extraordinary number of pre-main sequence (PMS) stars in the vicinity of the stellar association LH 95 in the Large Magellanic Cloud (LMC). Using the Advanced Camera for Surveys on-board the Hubble Space Telescope in wide-field mode we obtained deep high-resolution imaging of the main body of the association and of a nearby representative LMC background field. These observations allowed us to construct the color-magnitude diagram (CMD) of the association in unprecedented detail, and to decontaminate the CMD for the average LMC stellar population. The most significant result is the direct detection of a substantial population of PMS stars and their clustering properties with respect to the distribution of the higher mass members of the association. Although LH 95 represents a rather modest star forming region, our photometry, with a detection limit V ∼ 28 mag, reveals in its vicinity more than 2,500 PMS stars with masses down to ∼ 0.3 M_{⊙}. Our observations offer, thus, a new perspective of a typical LMC association: The stellar content of LH 95 is found to extend from bright OB stars to faint red PMS stars, suggesting a fully populated Initial Mass Function (IMF) from the massive blue giants down to the sub-solar mass regime.

Accepted for publication in Astrophysical Journal Letters Available from arXiv:0706.4377

A signature of the donor star in the extra-galactic X-ray binary LMC X-2

R. Cornelisse^{1,2}, D. Steeghs^{3,4}, J. Casares^{1}, P.A. Charles^{5,2}, I.C. Shih^{6,2}, R.I. Hynes^{7} and K. O’Brien^{8}

1Instituto de Astrofísica de Canarias, Spain 2University of Southampton, UK 3Harvard-Smithsonian Center for Astrophysics, USA 4University of Warwick, UK 5South African Astronomical Observatory, South Africa 6Michigan State University, USA 7Louisiana State University, USA 8European Southern Observatory, Chile

Two nights of phase-resolved medium resolution VLT spectroscopy of the extra-galactic low mass X-ray binary LMC X-2 have revealed a 0.32 ± 0.02 day spectroscopic period in the radial velocity curve of the He ii λ4686 emission line that we interpret as the orbital period. However, similar to previous findings, this radial velocity curve shows a longer term variation that is most likely due to the presence of a precessing accretion disk in LMC X-2. This is strengthened by He ii λ4686 Doppler maps that show a bright spot that is moving from night to night. Furthermore, we detect narrow emission lines in the Bowen region of LMC X-2,with a velocity of K_{em} = 351 ± 28 km s^{−1}, that we tentatively interpret as coming from the irradiated side of the donor star. Since K_{em} must be smaller than K_{2}, this leads to the first upper-limit on the mass function of LMC X-2 of f(M_{1}) ≥ 0.86 M_{⊙} (95% confidence), and the first constraints on its system parameters.

Accepted for publication in MNRAS Available from arXiv:0707.2018

A catalogue of low-mass X-ray binaries in the Galaxy, LMC, and SMC (4^{th} edition)

Q.Z. Liu^{1,2}, J. van Paradijs^{2} and E.P.J. van den Heuvel^{2}

1Purple Mountain Observatory, Chinese Academy of Sciences, Nanjing 210008, P.R. China 2Astronomical Institute ”Anton Pannekoek”, University of Amsterdam, Kruislaan 403, 1098 SJ Amsterdam, The Netherlands

We present a new edition of the catalogue of the low-mass X-ray binaries in the Galaxy and the Magellanic Clouds. The catalogue contains source name(s), coordinates, finding chart, X-ray luminosity, system parameters, and stellar parameters of the components and other characteristic properties of 187 low-mass X-ray binaries, together with a comprehensive selection of the relevant literature. The aim of this catalogue is to provide the reader with some basic information on the X-ray sources and their counterparts in other wavelength ranges (γ-rays, UV, optical, IR, and radio). Some sources, however, are only tentatively identified as low-mass X-ray binaries on the basis of their X-ray properties similar to the known low-mass X-ray binaries. Further identification in other wavelength bands is needed to finally determine the nature of these sources. In cases where there is some doubt about the low-mass nature of the X-ray binary this is mentioned. Literature published before 1 October 2006 has, as far as possible, been taken into account.

Published in A&A, 469, 807 (2007) Available from arXiv:0707.0544

High-mass X-ray binaries in the Magellanic Clouds

Q.Z. Liu^{1,2}, J. van Paradijs^{2} and E.P.J. van den Heuvel^{2}

1Purple Mountain Observatory, Chinese Academy of Sciences, Nanjing 210008, P.R. China 2Astronomical Institute ”Anton Pannekoek”, University of Amsterdam, Kruislaan 403, 1098 SJ Amsterdam, The Netherlands

We present a catalogue of high-mass X-ray binaries in the Magellanic Clouds. The catalogue lists source name(s), coordinates, apparent magnitudes, orbital parameters, and X-ray luminosity of 128 high-mass X-ray binaries, together with stellar parameters of the components, other characteristic properties and a comprehensive selection of the relevant literature. The aim of this catalogue is to provide easy access to the basic information on the X-ray sources and their counterparts in other wavelength ranges (UV, optical, IR, radio). Most of the sources are identified to be Be/Xray binaries. Some sources, however, are only tentatively identified as high-mass X-ray binaries on the basis of a transient character and/or a hard X-ray spectrum. Further identification in other wavelength bands is needed to finally determine the nature of these sources. In cases where there is some doubt about the high-mass nature of the X-ray binary this is mentioned. Literature published before 1 May 2005 has, as far as possible, been taken into account.

Published in A&A, 442, 1135 (2005)

The Magellanic Squall: Gas Replenishment from the Small to Large Magellanic Cloud

Kenji Bekki^{1} and Masashi Chiba^{2}

1School of Physics, University of New South Wales, Australia 2Astronomical Institute, Tohoku University, Sendai, Japan

We first show that a large amount of metal-poor gas is stripped from the Small Magellanic Cloud (SMC) and fallen into the Large Magellanic Cloud (LMC) during the tidal interaction between the SMC, the LMC, and the Galaxy over the last 2 Gyrs. We propose that this metal-poor gas can closely be associated with the origin of LMC’s young and intermediate-age stars and star clusters with distinctively low-metallicities with [Fe/H]< −0.6. We numerically investigate whether gas initially in the outer part of the SMC’s gas disk can be stripped during the LMC-SMC-Galaxy interaction and consequently can pass through the central region (R < 7.5 kpc) of the LMC. We find that about 0.7% and 18% of the SMC’s gas can pass through the central region of the LMC about 1.3 Gyr ago and 0.2 Gyr ago, respectively. The possible mean metallicity of the replenished gas from the SMC to LMC is about [Fe/H]= −0.9 to −1.0 for the two interacting phases. These results imply that the LMC can temporarily replenish gas supplies through the sporadic accretion and infall of metal-poor gas from the SMC. These furthermore imply that if these gas from the SMC can collide with gas in the LMC to form new stars in the LMC, the metallicities of the stars can be significantly lower than those of stars formed from gas initially within the LMC. We discuss these results in terms of the origin of distinctively low metallicities of an intermediate-age globular cluster (NGC 1718), very young star clusters (e.g. NGC 1984), and early-type stars in the inter-Cloud region close to the LMC.

Accepted for publication in MNRAS Letters Available from arXiv:0706.3780

On the Origin of the Kinematical Differences Between the Stellar Halo and the Old Globular Cluster System in the Large Magellanic Cloud

Kenji Bekki^{1}

1School of Physics, UNSW, Australia

We discuss structural and kinematical properties of the stellar halo and the old globular cluster system (GCS) in the Large Magellanic Cloud (LMC) based on numerical simulations of the LMC formation. We particularly discuss the observed possible GCS’s rotational kinematics (V/σ > 2) that appears to be significantly different from the stellar halo’s one with a large velocity dispersion (∼ 50 km s^{−1}). We consider that both halo field stars and old GCs can originate from low-mass subhalos virialized at high redshifts (z > 6). We investigate the final dynamical properties of the two old components in the LMC’s halo formed from merging of low-mass subhalos with field stars and GCs. We find that the GCS composed of old globular clusters (GCs) formed at high redshifts (z > 6) has little rotation (V/σ ∼ 0.4) and structure and kinematics similar to those of the stellar halo. This inconsistency between the simulated GCS’s kinematics and the observed one is found to be seen in models with different parameters. This inconsistency therefore implies that if old, metal-poor GCs in the LMC have rotational kinematics, they are highly unlikely to originate from the low-mass subhalos that formed the stellar halo. We thus discuss a scenario in which the stellar halo was formed from low-mass subhalos with no/few GCs whereas the GCS was formed at the very early epoch of the LMC’s disk formation via dissipative minor and major merging of gas-rich subhalos and gas infall. We suggest that there can be a threshold subhalo mass above which GCs can be formed within subhalos at high redshifts and thus that this threshold causes differences in physical properties between stellar halos and GCSs in less luminous galaxies like the LMC.

Accepted for publication in MNRAS Available from arXiv:0707.2188

Determination of the mass of the neutron star in SMC X-1, LMC X-4 and Cen X-3 with VLT/UVES

A. van der Meer^{1}, L. Kaper^{1,2}, M.H. van Kerkwijk^{3}, M.H.M. Heemskerk^{1} and E.P.J. van den Heuvel^{1,2}

1Astronomical Institute Anton Pannekoek, University of Amsterdam, The Netherlands 2Center for High-Energy Astrophysics, The Netherlands 3Department of Astronomy and Astrophysics, University of Toronto, Canada

We present the results of a spectroscopic monitoring campaign of the OB-star companions to the eclipsing X-ray pulsars SMC X-1, LMC X-4 and Cen X-3. High-resolution optical spectra obtained with UVES on the ESO Very Large Telescope are used to determine the radial-velocity orbit of the OB (super)giants with high precision. The excellent quality of the spectra provides the opportunity to measure the radial-velocity curve based on individual lines, and to study the effect of possible distortions of the line profiles due to e.g. X-ray heating on the derived radial-velocity amplitude. Several spectral lines show intrinsic variations with orbital phase. The magnitude of these variations depends on line strength, and thus provides a criterion to select lines that do not suffer from distortions. The undistorted lines show a larger radial-velocity amplitude than the distorted lines, consistent with model predictions. Application of our line-selection criteria results in a mean radial-velocity amplitude K_{opt} of 20.2 ± 1.1, 35.1 ± 1.5, and 27.5 ± 2.3 km s^{−1} (1σ errors), for the OB companion to SMC X-1, LMC X-4 and Cen X-3, respectively. Adding information on the projected rotational velocity of the OB companion (derived from our spectra), the duration of X-ray eclipse and orbital parameters of the X-ray pulsar (obtained from literature), we arrive at a neutron star mass of 1.06_{−0.10}^{+0.11}, 1.25_{−0.10}^{+0.11} and 1.34_{−0.14}^{+0.16} M_{⊙} for SMC X-1, LMC X-4 and Cen X-3, respectively. The mass of SMC X-1 is near the minimum mass (∼1 M_{⊙}) expected for a neutron star produced in a supernova. We discuss the implications of the measured mass distribution on the neutron-star formation mechanism, in relation to the evolutionary history of the massive binaries.

Accepted for publication in Astronomy and Astrophysics Available from arXiv:0707.2802

Hα emission from the Magellanic Bridge

Erik Muller^{1} and Quentin Parker^{2,3}

1Australia Telescope National Facility, CSIRO, Australia 2Department of Physics, Macquarie University, Australia 3Anglo-Australian Observatory, Australia

We present here a preliminary report and commentary of recently processed observations of Hα emission towards the Magellanic Bridge. These data have been analysed in an attempt to quantify the extent to which the stellar population is capable of reshaping the local ISM. We find that the Hα emission regions are small, weak and sparsely distributed, consistent with a relatively quiescent and inactive ISM where radiative and collisional ionisation is inefficient and sporadic. This suggests that energetic processes at the small scale (i.e. tens of pc) do not dominate the energy balance within the ISM of the Bridge, which therefore hosts a pristine turbulent structure, otherwise inaccessible within our own Galaxy. We find H emission that is well correlated with detected ^{12}CO(1-0) line emission (a proxy for molecular hydrogen), as well as other easily identified ring-like H i features.

Accepted for publication in Publications of the Astronomical Society of Australia. Available from arXiv:0706.4037

The Origin of Large-scale H i structures in the Magellanic Bridge

Erik Muller^{1} and Kenji Bekki^{2}

1Australia Telescope National Facility, CSIRO, Australia 2School of Physics, University of New South Wales, Australia

We investigate the formation of a number of key large-scale H i features in the ISM of the Magellanic Bridge using dissipationless numerical simulation techniques. This study comprises the first direct comparison between detailed H i maps of the Bridge and numerical simulations. We confirm that the SMC forms two tidal filaments: a near arm, which forms the connection between the SMC and LMC, and a counterarm. We show that the H i of the most dense part of the Bridge can become arranged into a bimodal configuration, and that the formation of a ”loop” of H i, located off the North-Eastern edge of the SMC can be reproduced simply as a projection of the counter-arm, and without invoking localised energy-deposition processes such as SNe or stellar winds.

Accepted for publication in MNRAS Letters Available from arXiv:0706.3544

Conference Papers

A Hubble View of Star Forming Regions in the Magellanic Clouds

Dimitrios A. Gouliermis^{1}, Thomas K. Henning^{1}, Wolfgang Brandner^{1}, Michael R. Rosa^{2}, Andrew E. Dolphin^{3}, Markus Schmalzl^{1}, Eva Hennekemper^{1}, Hans Zinnecker^{4}, Nino Panagia^{5}, You-Hua Chu^{6}, Bernhard Brandl^{7}, Sascha

P. Quanz^{1}, Massimo Robberto^{5}, Guido De Marchi^{8}, Robert A. Gruendl^{6} and Martino Romaniello^{2}

1Max-Planck-Institut für Astronomie, Königstuhl 17, 69117 Heidelberg, Germany 2European Southern Observatory, Karl-Schwarzschild-Straße 2, 85748 Garching bei München, Germany 3Raytheon Corporation, USA 4Astrophysikalisches Institut Potsdam, An der Sternwarte 16, 14482 Potsdam, Germany 5Space Telescope Science Institute, 3700 San Martin Drive, Baltimore, MD 21218, USA 6Department of Astronomy, University of Illinois, 1002 West Green Street, Urbana, IL 61801, USA 7Sterrewacht Leiden, P.O. Box 9513, Niels Bohrweg 2, 2300 RA Leiden, The Netherlands 8ESA, Space Science Department, Keplerlaan 1, 2200 AG Noordwijk, The Netherlands

The Magellanic Clouds (MCs) offer an outstanding variety of young stellar associations, in which large samples of low-mass stars (with sub-solar masses) currently in the act of formation can be resolved and explored sufficiently with the Hubble Space Telescope. These pre-main sequence (PMS) stars provide a unique snapshot of the star formation process, as it is being recorded for the last 20 Myr, and they give important information on the low-mass Initial Mass Function (IMF) of their host environments. We present the latest results from observations with the Advanced Camera for Surveys (ACS) of such star-forming regions in the MCs, and discuss the importance of Hubble for a comprehensive collection of substantial information on the most recent low-mass star formation and the low-mass IMF in the MCs.

Poster contribution, published in 41^{st} ELSAB Symposium: The Impact of HST on European Astronomy, 29 May to 1 June, 2007, ESTEC, Noordwijk, The Netherlands Available from arXiv:0706.0620

Dust-Gas Interaction in SNR 1987A

Eli Dwek^{1} and Richard G. Arendt^{2}

1Observational Cosmology Lab, NASA Goddard Space Flight Center, USA 2CRESST/UMBC, NASA Goddard Space Flight Center, USA

Multiwavelength observations of SNR1987A show that its morphology is rapidly changing at X-ray, radio, and optical wavelengths as the blast wave from the explosion expands into the circumstellar equatorial ring. Infrared emission arises from the interaction of dust grains with the hot X-ray emitting gas. We show that the IR emission provides important complementary information on the interaction of the SN blast wave with the circumstellar equatorial ring that cannot be obtained at any other wavelength.

Oral contribution, published in ”Supernova 1987A: 20 Years after - Supernovae and Gamma-Ray Bursters”, held in Aspen Colorado, USA, Feb 19-23, 2007. AIP Proceedings Available from arXiv:0705.3796

High Resolution Spectroscopy of SN 1987A’s Rings: He i 10830 and Hα from the Hotspots

Nathan Smith^{1} and Richard McCray^{2}

1Astronomy Department, University of California, 601 Campbell Hall, Berkeley, CA 94720, USA 2JILA, University of Colorado, 440 UCB, Boulder, CO 80309, USA

We present the first high-dispersion spectroscopy of He i 10830 from the hotspots in the ring around SN 1987A, obtained at Gemini South, spatially resolving the near and far sides of the ring. We compare these line profiles to similar echelle spectra of Hα and [N ii] 6583 obtained at the Magellan Observatory. We find that the He i profiles are much broader than H-alpha or [N II], but the He I profiles also have different shapes — they have enhanced emission at high speeds, with extra blueshifted emission on the north side of the ring, and extra redshifted emission on the south side. To explain this, we invoke a simple geometric picture where the extra He i emission traces hotter gas from faster shocks that strike the apex of the hotspots directly, while the Hα preferentially traces cooler lower-ionization gas from slower transverse shocks that penetrate into the sides of the ring.

Poster contribution, published in ”Supernova 1987A: 20 Years After: Supernovae and Gamma-Ray Bursters”, eds. S. Immler, K.W. Weiler & R. McCray. AIP, New York Available from arXiv:0705.3067

X-ray Evolution of SN 1987A

Bernd Aschenbach^{1}

1Max-Planck-Institut für extraterrestrische Physik, Giessenbachstraße, 85741 Garching, Germany

The X-ray observations of SN 1987A over the previous 20 years have seen the emergence of soft X-rays from the interaction of the explosion shock wave with the ambient medium. This shock wave is now interacting strongly with the inner ring and might have passed already the highest density regions. The emission can be described by thermal models with temperatures of ∼ 0.28 keV and ∼ 2.8_{−1.0}^{+0.7} keV, with perhaps some but little change over time. High resolution spectroscopy reveals a large variety of shock velocities ranging from a few hundred to many thousand km s^{−1}. Relative to the elemental abundances prevailing in the LMC the inner ring shows an overabundance of Si and S compared to the lighter elements and Fe, which suggests that the ring consists of highly processed matter dredged up in a binary merger event well before the explosion. The X-ray lightcurves between 0.5–2 keV and 3–10 keV differ significantly in slope, with the latter being much flatter but very similar to the radio light curve.

Oral contribution, published in ”Supernova 1987A: 20 Years after - Supernovae and Gamma-Ray Bursters” held in Aspen Colorado USA, Feb 19-23, 2007. AIP Proceedings Available from arXiv:0705.3406

HST’s view of the youngest massive stars in the Magellanic Clouds

M. Heydari-Malayeri ^{1}, M. R. Rosa^{2}, V. Charmandaris^{3}, L. Deharveng^{4}, F. Martins^{5}, F. Meynadier^{1}, D. Schaerer^{6,7} and H. Zinnecker^{8}

1LERMA, Paris Observatory, France 2ST-ECF, ESO, Germany; affiliated to the Space Telescope Operations Division, RSSD, ESA 3University of Crete, Greece 4Marseille Observatory, France 5MPE, Garching bei München 6Geneva Observatory, Switzerland; 7Observatoire Midi-Pyrénées, Laboratoire d‘Astrophysique, Toulouse, France 8Potsdam Astrophysical Institute, Germany

Accurate physical parameters of newborn massive stars are essential ingredients to shed light on their formation, which is still an unsolved problem. The rare class of compact H ii regions in the Magellanic Clouds (MCs), termed “high-excitation blobs” (HEBs), presents a unique opportunity to acquire this information. These objects (∼ 4′′ to 10′′, ∼ 1 to 3 pc, in diameter) harbor the youngest massive stars of the OB association/molecular cloud complexes in the MCs accessible through high-resolution near-IR and optical techniques. We present a brief overview of the results obtained with HST mainly on two HEBs, one in the LMC (N159-5) and the other in the SMC (N81).

Oral contribution, published in The 41^{st} ESLAB Symposium ”The Impact of HST on European Astronomy”, 29 May to 1 June 2007, ESTEC, Noordwijk, Netherlands; eds. G. De Marchi & D. Macchetto Available from arXiv:0707.1209

Job Adverts

Mayflower Scholarship for PhD position

There is a 4-year fully-funded PhD studentship available for any European student to work on the Magellanic Clouds from September 2007, or as soon as possible thereafter.

In particular, the programme is focussed on the evolved stellar populations of the MCs, especially Be/neutron star binary systems. The global aim is to understand why there is such a huge population of such objects in the SMC, as opposed to the LMC. The programme makes use of spaced-based X-ray data from RXTE & Chandra, as well as ground-based optical and IR observations. An interest in observing, data-analysis & data-interpretation is important.

The studentship (called a Mayflower Scholarship) is available to any European citizen. As part of the scholarship the student is expected to spend up to 20% of their time supporting the undergraduate teaching programme of the department. For further technical details on the studentship see: www.phys.soton.ac.uk/news/index.php?news=mayflower

For more information on the science, contact Malcolm Coe (mjcoe@soton.ac.uk).

See also http://www.astro.soton.ac.uk/∼mjc/res/

PdD opportunity on the VISTA Survey of the Magellanic Cloud System

The VMC — VISTA near-infrared survey of the Magellanic System (LMC, SMC, Bridge & Stream) — is one of the 6 public surveys approved by ESO to be conducted with the VIRCAM camera at the 4m VISTA telescope. It will cover 184 deg^{2} encompassing the D_{25} of both Clouds as well as major features delineated in the distribution of stars and H i gas, in the Y JK_{s} passbands. The survey depth — namely K_{s} = 20.3 at S/N = 10, i.e. 6 magnitudes deeper than 2MASS — has been chosen to reach the oldest main sequence turn-off across the entire system, and will allow us to determine the spatially-resolved star formation history (SFH) with unprecedented quality and to construct a detailed 3D map of the system.

A PhD scholarship is available to work at the Osservatorio Astronomico di Padova (OAPd), Italy, on the preparation and early data analysis of VMC. The successful candidate will follow the PhD School in Astronomy of the Padova University. The primary scope of the PhD project is in the realisation of artificial star tests and SFH-recovery tests, using both simulated data and the first VMC images (to be available in 2008), to be followed by the analysis of the RGB and AGB populations. The work is to be carried out together with Léo Girardi and Paola Marigo, from OAPd and the Astronomy Department respectively, and in collaboration with other VMC team members.

Interested candidates are encouraged to contact Léo Girardi at leo.girardi@oapd.inaf.it as soon as possible. The selection will be in Padova in November 2007, with the deadline for applications in October. The PhD school starts in January 2, 2008. The monthly net salary is of about 820 euro.

Knowledge of English and basic computational skills are required, previous experience with stellar photometry and/or studies of resolved stellar populations are assets.

See also http://stev.oapd.inaf.it/∼lgirardi/PhD project 2007.pdf